# GDP nowcasting with artificial neural networks: How much does long-term memory matter?

### 19 Dec 2023 v2

@@ 4.2 The nowcasting exercise
Our full dataset ranges from 1960:Q1 to 2022:Q4, so that it contains feature observations for every month and target observations for every third month (i.e., for every quarter). We assume that quarterly GDP measurements are available for months M3, M6, M9, and M12.161616FRED data releases associate these quarterly measurements with timestamps M1, M4, M7, and M10. However, using these timestamps would be highly counter-intuitive in this type of analysis. Following the literature, we assume that quarterly GDP measurements are published at the end of the current quarter. In reality, the first measurements are usually available only a few months after the end of the current quarter.
@@
where $\Omega_{v}$ represents the information set containing monthly regressor (indicator) series of length $l$ with measurements available until the $v$-th month. In line with the definitions of the different nowcasting scenarios, $v=\left[3t-2,\quad 3t-1,\quad 3t\right]$. Lastly, $\mathcal{M}$ refers to the underlying model. Based on Equation (7), we can also see that different nowcasting scenarios only differ in their corresponding information sets. More specifically, $v=3t-1$ refers to the 2-month nowcasting scenario. Thus, within a given quarter, training sequences of length $l$ include observations for monthly indicators until the second month. Intuitively, we expect to see the most accurate nowcasts in the 3-month nowcasting scenario ($v=3t$), where all intra-quarterly indicator data are available for prediction. In Equation (7), $\mathcal{M}$ denotes a concrete element of the set of our competitor ANNs (MLP, 1D CNN, RNN, LSTM, GRU). The underlying model determines what function $f()$ can be learned between the target variable and the regressors during the training (estimation) process. In the line of the above, the nowcast for quarter $q$ can be generated as follows:
@@
where $\mathbf{x}_{v}$ represents a row in the regressor vector, containing observations for $n$ number of features in month $v$. As Equation (8) suggests, we do not include any lagged values of the target variable in the regressor vector, which is denoted by $\varphi(\mathbf{x}_{v})$. Hence, if $l$ denotes the length of the training sequences (i.e. sequences of monthly indicators), the dimensionality of the regressor vector is $l\times n$. Accordingly, training sequences $\left(y_{q},\varphi(\mathbf{x}_{v})\right)$ are formed by pairing the actual target observations for quarter $q$ with the corresponding regressor vector. The training sequences are sampled then by a fixed-length rolling window where the size of the rolling window is determined by the sequence length ($l$). Since each (sub)sample of the features is related to a given quarterly target observation, successive rolling windows differ in values for their last 3 months. Consequently, they overlap in $l-3$ monthly values, provided that $l>3$.
@@
Figure 9 shows that the naive benchmark model fits the data surprisingly well during the first evaluation period, i.e., in periods of quasi-stable economic growth. However, it fails when large shocks (positive or negative) hit the economy, as that is apparent during the second test period. The economic impact of the COVID-19 pandemic resulted in a V-shaped recession: a sharp downturn in 2020:Q2 followed by a quick rebound in 2020:Q3. While signs of the economic downturn could have been seen already in the macro indicators for April, the naive benchmark model clearly fails to predict the forthcoming recession in Q2. Following the assumption of constant growth, it also misses the quick rebound in Q3, which affects its performance even worse. It is an interesting question: How accurately can our competitor models predict the V-shaped recession caused by the COVID-19 pandemic during the second evaluation period? We evaluate the nowcasting accuracy of our models in terms of relative root mean squared error (RMSE) and relative mean absolute error (MAE). As a benchmark, the constant growth model generates an RMSE of $0.490$ and an MAE of $0.361$ in the first test period. These values for the second test period are $2.980$ and $1.140$, respectively. These values form the basis of the model evaluation, and help indicate the competitor models’ absolute nowcasting accuracy. Table 1 summarizes the above considerations and points out the strong negative relationship between the volatility of GDP growth and the nowcasting performance of the benchmark model.
Table 1: Main descriptive statistics of GDP growth and the naive benchmark model’s nowcasting accuracy in the two evaluation periods.
@@
0.730 0.563 0.521
$\sigma\left(y_{t}\right)$
1.102 0.334 1.864
@@ 4.3 Results
In the following, we present the results of the empirical analysis. As the COVID-19 pandemic hit the US economy in 2020, we decided to evaluate our models on two different evaluation periods: The first evaluation period ranges from 2012:Q1 to 2019:Q4, so it does not reflect the economic impact of the COVID-19 pandemic. In comparison, the second evaluation period is from 2012:Q1 to 2022:Q4, containing those periods of the COVID crisis as well. We believe that the results from these two evaluation periods show a more balanced and well-rounded picture of the different models’ generalization capability. Clearly, these two test periods are very different in their characteristics. The first can be basically described as a period of balanced economic growth. Under such circumstances, the naive benchmark model is expected to be a strong contender for our competitor models. Consequently, competitor models should accurately predict even more intricate movements in economic growth to beat the benchmark model. In contrast, the second evaluation period would favor the models that can capture the sharp V-shaped recovery of 2020:Q2 and 2020:Q3. We believe that results from these two evaluation periods will provide a robust assessment about the generalization capability of the different competitor models, showing their strengths and weaknesses.
@@
As we proposed, 1D CNN represents a “sweet spot” in terms of architectural design between the simple time-agnostic MLP and the more complex gated recurrent networks. Based on the results of Table 2 and Table 3, the nowcasting performance of the 1D CNN seems very balanced as well, achieving the best accuracy for both evaluation periods.

### 29 Feb 2024 v3

@@ 4.2 The nowcasting exercise
Our full dataset ranges from 1960:Q1 to 2022:Q4, so that it contains feature observations for every month and target observations for every third month (i.e., for every quarter). We assume that quarterly GDP measurements are available for months M3, M6, M9, and M12.171717FRED data releases associate these quarterly measurements with timestamps M1, M4, M7, and M10. However, using these timestamps would be counter-intuitive in an empirical analysis focused on nowcasting. Following the literature, we assume that quarterly GDP measurements are published at the end of the current quarter. That corresponds with the reality since the first measurements (first estimates) are available about a month after the end of the current quarter.
@@
[…]
@@
where $\mathbf{x}_{v}$ represents a row in the regressor vector, containing observations for $n$ number of features in month $v$. As Equation (8) suggests, similarly to [40], we do not include any lagged values of the target variable in the regressor vector, which is denoted by $\varphi(\mathbf{x}_{v})$. Hence, if $l$ denotes the length of the training sequences (i.e. sequences of monthly indicators), the dimensionality of the regressor vector is $l\times n$. Accordingly, training sequences $\left(y_{q},\varphi(\mathbf{x}_{v})\right)$ are formed by pairing the actual target observations for quarter $q$ with the corresponding regressor vector. The training sequences are sampled then by a fixed-length rolling window where the size of the rolling window is determined by the sequence length ($l$). Since each (sub)sample of the features is related to a given quarterly target observation, successive rolling windows differ in values for their last 3 months. Consequently, they overlap in $l-3$ monthly values, provided that $l>3$.
@@
Figure 8 shows that the naive benchmark model fits the data surprisingly well during the first evaluation period, i.e., in periods of quasi-stable economic growth. However, it fails when large shocks (positive or negative) hit the economy, as that is apparent during the second test period. The economic impact of the COVID-19 pandemic resulted in a V-shaped recession: a sharp downturn in 2020:Q2 followed by a quick rebound in 2020:Q3. While signs of the economic downturn could have been seen already in the macro indicators for April, the naive benchmark model clearly fails to predict the forthcoming recession in Q2. Following the assumption of constant growth, it also misses the quick rebound in Q3, which affects its performance even worse. It is an interesting question: How accurately can our competitor models predict the V-shaped recession caused by the COVID-19 pandemic during the second evaluation period? We evaluate the nowcasting accuracy of our models in terms of relative root mean squared error (RMSE) and relative mean absolute error (MAE). As a benchmark, the constant growth model generates an RMSE of $0.513$ and an MAE of $0.389$ for the first test period. These measures for the second test period are $2.837$ and $1.110$, respectively. These values form the basis of the model evaluation, and also help to assess the absolute nowcasting accuracy of the competitor ANNs. Table 1 summarizes the above considerations and points out the strong negative relationship between the volatility of GDP growth and the nowcasting performance of the naive benchmark model.
Table 1: Main descriptive statistics of GDP growth and the nowcasting performance of the naive constant growth model in the two evaluation periods.
@@
0.736 0.612 0.555
$\sigma(y_{t})$
1.079 0.357 1.786
$100\cdot\left(\sigma(y_{t})/\bar{y}_{t}\right)$
146.554 58.312 321.667
@@ 4.3 Results
In the following, we present the results of the empirical analysis. As the COVID-19 pandemic hit the US economy in 2020, we decided to evaluate our models on two different evaluation periods: The first evaluation period ranges from 2012:Q1 to 2019:Q4, so it does not reflect the economic impact of the COVID-19 pandemic. In comparison, the second evaluation period is from 2012:Q1 to 2022:Q4, containing those periods of the COVID crisis as well. We believe that the results from these two evaluation periods show a more balanced and well-rounded picture of the different models’ generalization capability. Clearly, these two test periods are very different in their characteristics. The first can be basically described as a period of balanced economic growth. Under such circumstances, the naive benchmark model is expected to be a strong contender for our competitor models. Consequently, competitor models should accurately predict even the more intricate movements in economic growth to beat the benchmark model. In contrast, the second evaluation period would favor the models that can capture the sharp V-shaped recovery of 2020:Q2 and 2020:Q3. We believe that results from these two evaluation periods will provide a robust assessment about the generalization capability of the different competitor models, showing their strengths and weaknesses.
@@
As we proposed, the 1D CNN might represent a “sweet spot” in terms of architectural design between the simple time-agnostic MLP and the more complex (gated) recurrent networks. Based on the combined results of Table 2 and Table 3, the 1D CNN seems to be a very suitable ANN architecture for GDP nowcasting: It significantly outperforms the benchmark model in the first evaluation period while it generates the most accurate nowcasts for the second evaluation period.
@@ Appendix A Data appendix
(a) Real personal income (RPI): Original time series with trend.
(b) Real personal income (RPI): Log-differenced, detrended time series.
Figure A.1: Real personal income (RPI, ID: 1) before and after the corresponding data transformation (tcode: 5). Source: FRED-MD.